Melanie is 2 years older than her brother Tim. Ryan is 5 years younger than Melanie. If the sum of their ages is 56, how old is Melanie?
1
Expert's answer
2019-09-05T09:14:52-0400

Setting variables

 Melanie's age: X

Tim's age: Y

Ryan's age: Z


Melanie is 2 years older than her brother Tim.


"X=2+Y"


Tim's age expression


"Y=X-2" Equation (1)


Ryan is 5 years younger than Melanie.


"Z=X-5" Ecuation (2)


 the sum of their ages is 56


"X+Y+Z=56" Equation (3)


Replacing equation (1) and (2) in equation (3) and solving the equation for X


"X+X-2+X-5=56\\
3X-7=56 \\
3X=56+7\\
3X=63\\
X=\frac{63}{3}\\
X=21"


Melanie is "\boxed{21 \text{years old}}"